OpenStudy (anonymous):

8x-2/ 9 = 2/1 a. 9/4 b. 5/2 c. 9/2 d. 9/8

OpenStudy (mathlegend):

First, we know that 2/1 or anything over 1... is just itself, right?

OpenStudy (mathlegend):

So lets rewrite this as... 8x-2/9 = 2 Now, to get rid of the denominator of the fraction we need to multiply by that denominator on both sides.

OpenStudy (mathlegend):

\[9(\frac{ 8x-2 }{ 9 }) = 2(9)\]

OpenStudy (mathlegend):

The 9's on the left side cancel, right? On the right side it equals 18.

OpenStudy (mathlegend):

8x-2 = 18 Can you solve for x now @kira12341234 ?

OpenStudy (anonymous):

thank you so much i understand it much better now thank you :]

OpenStudy (mathlegend):

No problem :)